OpenStudy (haaileyyy):

Is this Correct? (-8ab+7a-5)-(-7ab+2) = -ab+7a+7

OpenStudy (rolldextro):

-8ab+7a-5+7ab-2=-ab+7a-7

OpenStudy (rolldextro):

Yes you are correct.

OpenStudy (mathstudent55):

No.

OpenStudy (mariamikayla):

(-8ab+7a-5)-(-7ab+2)= = -8ab+7a-5+7ab-2= = -ab+7a-2

OpenStudy (mathstudent55):

Do it one step at a time. Drop first set of parentheses. Distribute negative 1 through second set of parentheses. Collect like terms.

OpenStudy (mathstudent55):

@Mariamikayla Sorry, also incorrect.

OpenStudy (mathstudent55):

@rolldextro Your answer is different than the poster's, but you say she is correct. Do you mean your answer is incorrect?

OpenStudy (mathstudent55):

\((-8ab+7a-5)-(-7ab+2) =\) \(=-8ab+7a-5 +7ab-2 \) \(=-ab+7a-7 \)

OpenStudy (mariamikayla):

@mathstudent55 aah, yes, sorry... it was -ab+7a-7 ...i've forgot to add -5

OpenStudy (mathstudent55):

@Haaileyyy Your only mistake was with the constant terms. What is -5- 2?

OpenStudy (haaileyyy):

this is how i did it \[(-8ab+7a-5)-(-7ab+2)\] \[-8ab+7a-5+7ab-2\] \[-ab+7a+7\]

OpenStudy (haaileyyy):

oohhh its supposed to be -7 and not positive

OpenStudy (mathstudent55):

Your second line is correct. -5 - 2 = -7

OpenStudy (mathstudent55):

You got it now.

OpenStudy (haaileyyy):

Thank you so much

OpenStudy (mathstudent55):

yw